Please Help Me!!!!
Suppose instead that the town experiences quadratic growth of the form P= 7t^2=Po where t is the time of years from 2010. The Po=1000
A)Insert the value of Po that your group has decided upon into the equation. Use t^2 to type t-squared
B) Fill in the following chart.
Year(t) Population(p)
T=0(2010) Po=
T=1(2011) Po=
T=2(2012) Po=
T=3(2013) Po=
T=6(2016) Po=
C) Use your equation from part a to approximate how may years it will take for the population to reach 12,000. Round to the nearest whole year when necessary.

Answer by RAY100(1637)

Please check to see that this is your problem
.
P =7t^2 +c
.
P(0) = 1000=7t^2 +c = 0 +c,,,,or c=1000
.
P(0) = 7t^2 + 1000= 1000
P(1) = 7*1^2 +1000= 1007
P(2) = 7*2^2 +1000= 1028
P(3) = 7*3^2 +1000= 1063
P(4) = 7*4^2 +1000= 1112
P(5) = 7*5^2 +1000= 1175
p(6) = 7*6^2 +1000= 1252
.
12,000 = 7*t^2 +1000
11,000 = 7*t^2
1571.4 = t^2
39.64 = t